OpenStudy (anonymous):

Find the sum to n terms of the series 0.6+0.66+0.666+.....................

OpenStudy (anonymous):

oh no sorry

OpenStudy (anonymous):

S = 0.6 + 0.66 + 0.666 + 0.6666........ S - S = 0.6 + 0.06 + 0.006 + 0.0006 ......- a n. an = 0.6 ( 10 ^-n - 1)/ (-0.9) An = 2/3 ( 1 - 10^-n) = 2/3 -2*10^-n/3 Now try using this general term.

OpenStudy (ash2326):

We have \[0.6+0.66+0.666+0.6666\] Let's take 6 out from the series \[6(0.1+0.11+0.111+0.1111+....\] Now multiply and divide by 9 \[\frac{6}{9}(0.9+0.99+0.999+0.9999+....\] \[\frac{6}{9}(1-0.1+1-0.01+1-0.001+1-0.0001+....\] \[\frac{6}{9}(1-10^{-1}+1-10^{-2}+1-10^{-3}+...\] all 1s add up to n \[\frac{6}{9}(n-\frac{1}{10}\times \frac{1-(\frac{1}{10})^{n+1}}{1-\frac{1}{10}}\]

OpenStudy (zarkon):

should you have \[\frac{6}{9}\left(n-\frac{1}{10}\times \frac{1-(\frac{1}{10})^{n}}{1-\frac{1}{10}}\right)\]

OpenStudy (anonymous):

I got the same result. And yes that's correct. Did you not understand my method? I ould explain it again. ash2326 ---> Nice. :D